### v1

# references.bib

ABSTRACT
OBJECTIVE: Ensuring that machine learning (ML) algorithms are safe and effective within all patient groups, and do not disadvantage particular patients, is essential to clinical decision making and preventing the reinforcement of existing healthcare inequities. The objective of this tutorial is to introduce the medical informatics community to the common notions of fairness within ML, focusing on clinical applications and implementation in practice.
TARGET AUDIENCE: As gaps in fairness arise in a variety of healthcare applications, this tutorial is designed to provide an understanding of fairness, without assuming prior knowledge, to researchers and clinicians who make use of modern clinical data.
SCOPE: We describe the fundamental concepts and methods used to define fairness in ML, including an overview of why models in healthcare may be unfair, a summary and comparison of the metrics used to quantify fairness, and a discussion of some ongoing research. We illustrate some of the fairness methods introduced through a case study of mortality prediction in a publicly available electronic health record dataset. Finally, we provide a user-friendly R package for comprehensive group fairness evaluation, enabling researchers and clinicians to assess fairness in their own ML work.
INTRODUCTION
There are myriad potential applications of machine learning (ML) in healthcare, including automated disease detection, computed-aided diagnosis, and personalized treatment planning [rajpurkar2022ai]. However, there is substantial evidence that, without appropriate forethought and planning, ML models can exacerbate existing inequities in healthcare by making less accurate decisions for certain groups or individuals [grote_enabling_2022]. Within medical imaging, state of the art ML models used for disease diagnosis, risk prediction, and triage management are known to underperform within minority groups defined by protected attributes, including sex, race, and ethnicity [rajpurkar_chexnet_2017, ting_development_2017, liu_comparison_2019, lee_machine_2020, seyyed-kalantari_chexclusion_2020, larrazabal_gender_2020, seyyed-kalantari_underdiagnosis_2021, puyol-anton_fairness_2022, bressem_deep_2021, roest_comparative_2022, lin_evaluate_2023]. A recent study found that deep learning models used to detect 14 common diseases from chest X-rays significantly under-diagnosed Hispanic females, potentially resulting in delays in treatment if deployed in practice [seyyed-kalantari_underdiagnosis_2021]. In the same vein, ML models developed with electronic health record (EHR) and claims data exhibit similarly biased performance by under-performing in particular subgroups [li_evaluating_2023, banda_characterizing_2023]. In a landmark paper, researchers found that a widely-used commercial risk prediction model for identifying patients with complex health needs for high-risk care management performed significantly worse for Black patients relative to white patients [obermeyer_dissecting_2019]. Consequently, the model reduced the number of black patients designated for necessary extra care by over 50%.
As applications of ML in healthcare become commonplace, it is crucial to recognize, account for, and mitigate such disparities in model performance. Broadly, an ML model is said to be fair if it does not discriminate against an individual or group [barocas_fairness_2023]. Concepts of fairness have been extensively studied across various disciplines, including social choice theory, game theory, economics, and law [young_equity_1995, roemer_equality_1998, roemer_theories_1998, rawls_theory_1971, rawls_justice_2001, hutchinson201950, dwork_fairness_2012]. Building on these principles, the subfield of fairness within ML provides a framework for evaluating and mitigating bias throughout the model development and deployment process.
Although fairness has been an extremely active area of research over the past decade, evaluating and implementing fairness within ML for healthcare is relatively nascent [wang_bias_2022, feng_fair_2022, liu_translational_2023]. For instance, a recent systematic review of articles utilizing ML for EHR-based phenotyping found that only 5% of studies assessed fairness [yang_machine_2023]. Similarly, a review of EHR-based prediction models found that most studies only investigate the overall performance of ML models and do not interrogate potential biases [chen2024unmasking]. Beyond EHR applications, several scoping reviews of clinical ML models found that the adoption of fairness remains inconsistent, which can be partially attributed to a knowledge gap between ML and clinical researchers [liu_translational_2023, plecko_causal_2022, anderson2024algorithmic]. Moreover, more than two dozen definitions of fairness have been proposed, making it a particularly challenging domain to navigate.
To increase familiarity with this important topic within medical informatics, we provide a didactic introduction to the fundamental concepts and methods of fairness within ML. This tutorial is designed to serve as a starting point for informaticians and clinical researchers exploring the rich and nuanced fairness literature and is complementary to book chapters, workshops, articles, and software that are references on the topic. Supplementary Tables S1 and S2 provide a detailed comparison of our tutorial with existing literature and software. In contrast to other references, we focus specifically on healthcare applications and include a fully reproducible case study of mortality prediction on publicly available EHR data from the Medical Information Mart for Intensive Care (MIMIC) clinical database [raffa2016clinical]. We also introduce a user-friendly R package that offers a full suite of methods for group fairness evaluation, some of which are not currently available in existing software.
We begin our tutorial with a discussion of why ML models can be unfair by introducing biases that can arise throughout the model development process using examples from the literature. We then provide a formal introduction to the most common fairness framework, group fairness, and illustrate how many group fairness definitions are incompatible, meaning that they cannot be simultaneously satisfied. We also present a case study on mortality prediction using our group fairness R package to reinforce these concepts. We then introduce individual and causal fairness and close with a discussion of existing approaches and some ongoing research.
WHY ML MODELS ARE UNFAIR
ML models are unfair due to biases in the data, in the model, and/or in the deployment of the model (Figure 1) [rajkomar_ensuring_2018, mehrabi_survey_2021, fazelpour_algorithmic_2021, mhasawade2021machine, wang_bias_2022, xu_algorithmic_2022, drukker_toward_2023, ferrara_fairness_2024]. The data used in healthcare applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [iliadis2016critical, suresh2021framework]. In addition to societal bias, data can also fail to represent the population of interest due the sampling method, time of collection, or data quality issues (e.g., missing, mismeasured, or not enough data). In the fairness literature, data with these undesirable properties are informally referred to as “biased” [chouldechova_fair_2017, barocas_fairness_2023]. Choices made during model training and evaluation can also incorporate and amplify bias in the data, such as selecting an inappropriate outcome, model, or metric(s) and data used for evaluation. During deployment, bias can be introduced when users selectively disregard or overly trust a model’s outputs, which can further reinforce existing societal bias. Table 1 details some common sources of bias together with an example from the healthcare literature. A more comprehensive review of causes of unfairness and applications outside of healthcare can be found in [suresh2021framework] an [mehrabi_survey_2021].
It is important to note bias should be a primary consideration throughout the ML pipeline, from problem formulation to model deployment [gichoya2023ai]. Building on existing AI reporting guidelines, a bias evaluation checklist has recently been proposed to enable practitioners to systematically and holistically address bias in clinical predictive models [wolff2019probast, liu2020reporting, wang_bias_2022]. The fairness criteria we introduce in the subsequent section play an important role in this process. They are typically integrated directly into model training to ensure fairness or used as evaluative metrics to measure disparities after the model has been trained [pessach_review_2022].
Figure 1: Sources of bias in ML. Unfairness arises from bias in the process of generating data, modeling data, and deploying the model. This figure is adapted from [mitchell_algorithmic_2021] and [mbakwe2023fairness].
(a) Bias in the data
Type of Bias Definition Example
Historical Bias The data reflects long-standing societal disparities encoded within the data over time. Clinical word embeddings trained on large corpora of text, such as clinical notes from healthcare systems, reflect biases about ethnic minorities [zhang2020hurtful].
Selection Bias The data is not representative of the population of interest. Data in volunteer-based biobanks does not reflect the general population as participation rates are higher among women, older age groups, and individuals with higher socioeconomic status. [fry_comparison_2017].
Measurement Bias The data contains variables that are collected or measured inaccurately. Gender identity is often recorded in EHRs without provider education and therefore does not accurately represent an individual’s identity [kronk_transgender_2022].
Temporal Bias The data captures a specific time period that may not reflect current or future conditions. Pediatric mental health visits were lower than expected for individuals with lower socioeconomic status during the first year of the COVID-19 pandemic [toulany_sociodemographic_2023].
Minority Bias The data lacks adequate representation from the minority group for the model to accurately learn. Most of the data for genetic studies is from European ancestry populations [martin_clinical_2019].
Missing Data Bias The data has variables that are incompletely measured. Patients from low-income backgrounds have higher rates of missing medical measurements, which is often due to inequities in access to healthcare [Chapman_Exploring_2022].
(b) Bias in the model
Type of Bias Definition Example
Label Bias An imperfect proxy is selected to train a model instead of the outcome of interest. Healthcare cost was used as a proxy for healthcare need in a commercial algorithm used to identify patients for high-risk care management programs. The algorithm significantly under-identified black patients as less money is spent on black patients relative to similarly healthy white patients [obermeyer_dissecting_2019].
Algorithmic Bias Properties of a model and/or its training algorithm create or amplify bias in the data. Differentially private prediction models for mortality performed poorly for minority groups [suriyakumar2021chasing].
Evaluation Bias An inappropriate choice of benchmark data or metrics are used for evaluation. EuroSCORE, a widely used cardiac surgery risk evaluation model, was only evaluated on European patients and exhibited decreased performance in other populations [debray_guide_2017].
(c) Bias in deployment
Type of Bias Definition Example
Automation Bias Humans overly trust model outputs, sometimes even against their own knowledge. A computer-aided detection system that scans mammograms and marks suspicious areas of potential cancer features had lower sensitivity for women aged 40–49 compared to older age groups. A clinician using this system can disproportionately miss cancers for this age group [lehman_diagnostic_2015].
Dismissal Bias Humans ignore model recommendations, often due to frequent false alerts. If the computer-aided detection system also has a lower positive predictive value for women aged 40–49, then clinicians may disregard its recommendations, as they are more likely to be false positives [ancker2017effects].
Table 1: Common biases that arise in machine learning (ML) applications in healthcare.
WHAT IT MEANS FOR A MODEL TO BE FAIR
In spite of the numerous fairness definitions, existing criteria primarily fall into three categories: group fairness, individual fairness, and causal fairness. Group fairness is a popular framework which includes criteria that deem a model fair if its predictions are similarly accurate or calibrated across a predefined set of groups. These groups are most often defined by a protected attribute(s) such as age or race [^1] Sensitive attribute and protected attribute are used interchangeably [caton_fairness_2023]. Fairness is also evaluated across groups defined by social determinants of health (e.g., income and education) [mccradden2020ethical]. [mehrabi_survey_2021]. Common protected attributes used in the healthcare literature are summarized in Table 2. In contrast, individual fairness requires that the model provide similar predictions to similar individuals based on user-defined similarity metrics [dwork_fairness_2012]. Causal fairness criteria rely on causal models to link observed disparities in model performance to their underlying cause [plecko2022causal]. Group fairness criteria are referred to as “oblivious” as they equate fairness with parities based solely on the distribution of the data, while individual and causal fairness criteria require additional context in the form of the similarity metrics and causal models, respectively [hardt_equality_2016, mitchell_algorithmic_2021]. We therefore only illustrate group fairness in our case study and refer readers to several excellent resources for individual and causal fairness. As we will soon see, many fairness criteria are incompatible in that they cannot be simultaneously satisfied. A taxonomy of fairness, including some different types of group, individual, and causal fairness and their incompatibilities, is presented in Figure 2.
Attribute Availability in MIMIC-II Examples
Age [seyyed-kalantari_chexclusion_2020, garriga_machine_2022, schuch_fairness_2023, yfantidou_accuracy_2023]
Disability [garriga_machine_2022]
Marital Status [wang_comparison_2022]
National Origin [yfantidou_accuracy_2023]
Race [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, garriga_machine_2022, schuch_fairness_2023, yuan_algorithmic_2023, yfantidou_accuracy_2023, rajkomar_ensuring_2018, wang_bias_2022]
Sex [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, garriga_machine_2022, yuan_algorithmic_2023, yfantidou_accuracy_2023]
Socioeconomic Status [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, schuch_fairness_2023, yfantidou_accuracy_2023, rajkomar_ensuring_2018, wang_bias_2022]
Table 2: Commonly used protected attributes in ML in healthcare. Examples of studies that evaluate fairness with respect to each protected attribute and the availability of each attribute in the the MIMIC-II database used in our case study.
Figure 2: Taxonomy of fairness criteria. Fairness criteria primarily fall into three categories: group fairness, individual fairness, and causal fairness. In contrast to group fairness criteria which can be inferred from the distribution of the data, individual and causal require additional considerations based on the context. Many fairness criteria cannot be simultaneously satisfied and incompatibilities are depicted with a red X.
NOTATION
We introduce a small amount of notation to introduce mathematical definitions of various criteria that complement our discussion in the main text. We denote the outcome of interest as $Y$, the features used for model training as ${\bf X}$ [^2] Note that ${\bf X}$ does not have to be a vector; it could be an image, text, tensor, or any other type of data., and the categorical variable for the protected attribute as $A$. ${\bf X}$ may or may not contain $A$ and we discuss this issue when we introduce individual fairness. We let $S=\hat{f}({\bf X})$ be the output from the ML model, where $\hat{f}$ is learned from a set of training data. For example, when $Y$ is continuous, $S$ is simply the predicted value of $Y$. In the context of classification, $Y$ is a binary label for membership in the positive or negative class and $S$ is the predicted probability of being in the positive class, referred to as the risk score. The final prediction, $D$, is positive if $S$ exceeds a pre-determined classification threshold and negative otherwise.
GROUP FAIRNESS
Group fairness criteria require ML models to perform similarly across groups of interest and are the most popular fairness framework [awasthi2020beyond]. The criteria primarily fall into three categories: independence, separation, and sufficiency [berk_fairness_2018, barocas_fairness_2023]. We provide mathematical definitions of various criteria, their interpretation in the context of a healthcare example, and include additional metrics that do not fall within the three categories in Table 3. In the following discussion of group fairness, we focus on classification as it is most widely used and well-studied [mitchell_algorithmic_2021, mehrabi_survey_2021, castelnovo_clarification_2022]. Supplementary Section 2.1 reviews measures of discrimination and calibration as many group fairness definitions for classification are based on them [mehrabi_survey_2021]. Additional approaches for settings with continuous outcomes, continuous and categorical protected attributes that define more than two groups, and multiple protected attributes (i.e., subgroup fairness) are discussed in Supplementary Sections 2.2 – 2.5.
INDEPENDENCE
[…]
SEPARATION
Separation requires that the risk score does not depend on the protected attribute within the positive and negative classes (i.e., $S\perp\!\!\!\perp A\mid Y)$. This implies that, among individuals in the positive (or negative) class, the rate of making a positive (or negative) prediction is consistent across groups. Common separation-based metrics therefore aim to equalize error rates across the groups, including the false negative rate (known as equality of opportunity), false positive rate (known as predictive equality), or both (known as equalized odds). These metrics have been widely used in clinical settings as misclassified predictions can have severe consequences in practice. For instance, a recent study evaluated equality of opportunity in the context of chest X-ray diagnostics, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [seyyed-kalantari_underdiagnosis_2021]. State-of-the-art computer vision models were found to have higher false negative rates in several intersectional under-served subpopulations, such as Hispanic female patients. There are also situations where prioritizing equalized odds is appropriate [hardt_equality_2016]. In the context of rapid screening tools for COVID-19 in emergency departments, a high false negative rate can lead to insufficient monitoring of COVID-19 while a high false positive rate can result in unnecessary additional testing [yang2023algorithmic]. Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [kleinberg2017inherent].
SUFFICIENCY
Sufficiency aims to equalize error rates among individuals with similar predictions [castelnovo_clarification_2022]. Formally, sufficiency requires that the label does not depend on the protected attribute given the risk score ( i.e., $Y\perp\!\!\!\perp A\mid S)$. The risk score is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023]. Common sufficiency-based metrics focus on equalizing the positive predictive value (known as predictive parity), both the positive and negative predictive value (known as conditional use accuracy equality), and calibration (known as well-calibration )[chouldechova_fair_2017, berk_fairness_2018, kleinberg2017inherent, guo_calibration_2017]. Within healthcare, ML models that fail to meet sufficiency criteria have led to inequitable resource allocation. For example, in models predicting in-hospital patient deterioration, achieving predictive parity ensures that the allocation of intensive care unit beds is based solely on medical need rather than protected attributes so that care is provided to those most in need [byrd2023validation]. In terms of calibration-based metrics, a commercial algorithm to enroll US patients into high-risk was shown to violate well-calibration. At equivalent risk scores, Black patients were in significantly poorer health than white patients and therefore had to be sicker for enrollment into the program [obermeyer_dissecting_2019].
INCOMPATIBILITIES
Independence, sufficiency, and separation provide different perspectives on what it means for a model to be fair. Except under highly restrictive conditions, it is not possible for an algorithm to fulfill all criteria at the same time [hardt_equality_2016, chouldechova_fair_2017, corbett-davies_measure_2023]. As such, it is important for researchers to choose which group fairness considerations are most relevant to the specific context.
More specifically, the following pairs are incompatible in the sense that they cannot generally be simultaneously satisfied: independence and sufficiency, independence and separation, and separation and sufficiency. A basic requirement for any of these pairs to be satisfied is that the outcome and the protected attribute are marginally independent (i.e., $Y\perp\!\!\!\perp A$). In classification problems, this implies that the probability of being in the positive class is the same across groups, which can be violated in many clinical applications (e.g., disease prevalence varies by age). Our case-study in the next section reinforces these incompatibilities and additional mathematical details are provided in Supplementary Section 3.
Metric Definition Interpretation Relax-
ation
Independence ($S\perp\!\!\!\perp A$)
Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In all-cause mortality prediction, the probability of the model predicting mortality is the same for males and females.
Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In all-cause mortality prediction, the probability of the model predicting mortality is the same for males and females after adjusting for age and pre-existing conditions.
Separation ($S\perp\!\!\!\perp A\mid Y$)
Equalized Odds $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\text{and}$
$P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$
In predicting asthma exacerbation in children, the rates of false positives (children without exacerbation incorrectly identified as having exacerbation) and false negatives (children with exacerbation incorrectly identified as not having exacerbation) are the same for children from low and high socio-economic classes [juhn2022assessing].
Predictive Equality $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting cardiovascular disease (CVD) diagnoses CVD in males and females without CVD at the same rate [Vargae002395].
Equal Opportunity $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting a suspicious finding from a chest X-rays misses findings in males and females at the same rate [ueda_artificial_2021].
Balance for Positive Class $\mathbb{E}(S\mid Y=1,A=a_{0})=\mathbb{E}(S\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average predicted score of people with lung cancer is equal for males and females [ueda_artificial_2021].
Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average risk scores among males and females without lung cancer are equal [ueda_artificial_2021].
Sufficiency ($Y\perp\!\!\!\perp A\mid S$)
Conditional Use Accuracy Equality $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})$ and
$P(Y=0\mid D=0,A=a_{0})=P(Y=0\mid D=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$
In predicting hospital readmission, the probability of being readmitted given the model makes that prediction and the probability of not being readmitted given the model makes that prediction is the same for Black and white patients [raza2023auditing].
Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that prediction is the same for Black and white patients [raza2023auditing].
Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all classification thresholds for males and females [roosli_2022_peeking].
Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all classification thresholds [roosli_2022_peeking].
Other
Brier Score Parity $\mathbb{E}[(Y-S)^{2}\mid A=a_{0}]=\mathbb{E}[(Y-S)^{2}\mid A=a_{1}],\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to enroll patients into high-risk care management programs, the mean squared error between the risk score and the label is the same for Black and White patients [obermeyer_dissecting_2019].
Overall Accuracy Equality $P(D=Y\mid A=a_{0})=P(D=Y\mid A=a_{1})\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the probability of a correct classification (i.e., correctly classifying a patient as having or not having lung cancer) is the same for males and females [ueda_artificial_2021].
Treatment Equality #$\frac{\#\{D=0,Y=1,A=a_{0}\}}{\#\{D=1,Y=0,A=a_{0}\}}=\frac{\#\{D=0,Y=1,A=a_{1}\}}{\#\{D=1,Y=0,A=a_{1}\}}\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the ratio of the false negatives and false positives is the same for both males and females [ueda_artificial_2021].
Table 3: Common group fairness criteria. Mathematical definitions of group fairness criteria, an interpretation in the context of an example from the literature, and an indication of whether the criterion is a relaxation of independence, separation, or sufficiency. Symbols: #=#absent\#=# = number of, $P=$ probability, $E=$ expected value. Notations: $Y:$ outcome, ${\bf X}:$ features used for model training, $A:$ protected attribute that takes value in the set $\mathcal{A}$, $S:$ risk score, $D:$ final prediction based on thresholding $S$, ${\bf Z}$: additional set of features.
CASE STUDY
As group fairness metrics are often used to evaluate bias after model training, we illustrate their application in a case study using a publicly available dataset on 1,776 ICU patients from MIMIC-II [mitcriticaldata2016secondary, raffa2016clinical]. We focus on predicting 28-day mortality and evaluate disparities in model performance across the protected attribute of sex which is recorded as either male or female. [^4] MIMIC-IV is the most current release of data, but publicly available data exists for MIMIC-II. We also acknowledge that sex is not binary, but recorded this way in the MIMIC-II data [lancaster2023complexities]. Our purpose is to highlight the capabilities of our R package and reinforce the concepts of independence, separation, and sufficiency. We therefore evaluate all metrics rather than selecting one for the problem at hand. Table 4 shows point estimates and 95% confidence intervals (CIs) for both the difference and ratio of various metrics, which are easily obtained from our package. Details of the data processing and modeling are in Supplementary Section 4 and the methods used for constructing the 95% CIs are in Supplementary Section 5. All analyses can be replicated using the code and tutorial at: https://jianhuig.github.io/FairnessTutorial/index.html.
The group fairness incompatibilities are illustrated in our case study. Independence is likely violated as the estimated mortality rates are for males and females are different (14% vs. 19%, respectively). Our model only satisfies sufficiency criteria, but neither independence nor separation criteria. The measures of statistical parity indicate that the model predicts a significantly higher mortality rate for females, even after conditioning on age. Given that the estimated mortality rates are different in females and males, strictly enforcing independence may not be advisable. With respect to separation, we observe that equal opportunity is not satisfied. Practically, these disparities could lead to higher rates of undetected mortality risk among male individuals and insufficient care. On the other hand, the sufficiency criterion of predictive parity is satisfied. Lastly, three additional metrics from Table 3 that do not fall within the three aforementioned categories that measure calibration and/or discrimination are satisfied.
Fairness Metric Female Male Difference 95% CI for the difference Ratio 95% CI for the ratio
Independence
Statistical Parity 17% 8% 9% [5%, 13%] 2.12 [1.49, 3.04]
Conditional Statistical Parity (Age $\geq 60$) 34% 21% 13% [5%, 21%] 1.62 [1.17, 2.23]
Separation
Equal Opportunity 38% 62% -24% [-39%, -9%] 0.61 [0.44, 0.86]
Predictive Equality 8% 3% 5% [2%, 8%] 2.67 [1.39, 5.12]
Balance for Positive Class 46% 37% 9% [4%, 14%] 1.24 [1.09, 1.42]
Balance for Negative Class 15% 10% 5% [3%, 7%] 1.5 [1.28, 1.75]
Sufficiency
Predictive Parity 62% 66% -4% [-21%, 13%] 0.94 [0.72, 1.23]
Other
Brier Score Parity 9% 8% 1% [-1%, 3%] 1.12 [0.89, 1.42]
Overall Accuracy Equality 87% 88% -1% [-5%, 3%] 0.99 [0.94, 1.04]
Treatment Equality 5.11 13.6 -8.49 [-32.96, 15.98] 0.38 [0.1, 1.34]
Table 4: Group fairness evaluation in the MIMIC-II mortality case study. Each metric quantifies disparities in model performance across males and females. Differences and ratios are presented alongside 95% confidence intervals (CIs).
INDIVIDUAL FAIRNESS
In contrast to group fairness criteria that target the average performance of a model across groups, individual fairness ensures that “any two individuals who are similar with respect to a particular task should be classified similarly” [dwork_fairness_2012]. One of the first works on individual fairness, fairness through awareness (FTA), formalized the notion of similarity by defining metrics that measure the distance between individuals and their predictions. Roughly speaking, fairness is achieved if the distance between the prediction for any two individuals is at most the distance between them. The choice of distance metric for the predictions depends on the type of outcome under study. For example, the Euclidean distance is used for logit transformed predicted probabilities in classification while the absolute difference of the hazard function has been used for survival analysis [keya2021equitable, rahman2022fair]. However, the metric used to measure the similarity between individuals has been a focus of ongoing research as it encodes exactly how individuals are similar, which relies on an “awareness” of the context. In the original FTA paper, the authors suggest leaving the choice of metric to domain experts. As there is currently no agreed upon standard, various strategies have been proposed to learn the metric from available data [kim2018fairness, ilvento2019metric, mukherjee2020two]. A recent scoping review on individual fairness in healthcare provides a comprehensive list of individual fairness methods and corresponding software [anderson2024algorithmic].
It is important to note that FTA is a counter approach to fairness through unawareness (FTU), which is intended to be a catch all solution to prevent bias by not explicitly including protected attributes into modeling [kusner_counterfactual_2017]. However, removing the protected attributes from a model warrants careful consideration. First, protected attributes may serve as critical predictors for the outcome of interest, such as age in sepsis or cardiovascular disease prediction and race in cancer screening models [damen2016prediction, chicco2020survival, cirillo2020sex]. Excluding these predictors can diminish the overall predictive accuracy and impact all individuals adversely or even lead to bias against the majority group. Second, protected attributes are often highly correlated with non-protected attributes. Simply removing protected attributes from the model does not prevent the model from inferring them from other attributes. FTU is therefore not always ethical or achievable [banja2023fairness]. Lastly, individual fairness is motivated by an inherent weakness in group fairness criteria that only consider average model performance within groups. That is, there are situations in which group fairness can be satisfied, but individuals within a group can be discriminated against. The incompatibilities among individual and group fairness have been examined in [friedler_impossibility_2016, binns_apparent_2020, castelnovo_clarification_2022, xu_compatibility_2024]. For example, when group-specific prevalences of an outcome differ, the predictions must be adjusted to achieve statistical parity and obtain an equal rate of positive predictions across the groups. This can cause individuals with similar characteristics to receive different predictions based solely on their protected attributes, which contradicts the principle of individual fairness.
CAUSAL FAIRNESS
We conclude with a brief overview of causal fairness, a topic that has gained popularity in recent years and itself worth a tutorial in the context of healthcare. As the name suggests, causal fairness criteria evaluate the causal relationship between a protected attribute and a model’s prediction rather than a performance measure or distance metric as in group and individual fairness [plecko_causal_2022]. A deep understanding of causal fairness requires knowledge of causal inference and we refer readers to several general and fairness-focused reviews on the topic [pearl2009causal, hernan2010causal, loftus2018causal, fawkes2022selection, makhlouf2022survey, carey2022causal, nilforoshan_causal_2022]. We introduce one way of incorporating causality into what it means to be fair based on counterfactuals, which are simply “what-if” statements [pearl2009causal, hernan2010causal]. In the context of fairness, the counterfactual is most often the unobserved prediction that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new]. One of the first counterfactual-based fairness criteria, simply called counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables are held constant [kusner_counterfactual_2017].
To better explain this concept, we briefly introduce directed acyclic graphs (DAGs). DAGs consist of nodes that represent variables and edges that connect them to illustrate their causal relationships and are used to formulate many causal fairness criteria [pearl2009causal]. For purposes of illustration, consider an over-simplified view of the world in predicting length of stay in the hospital. The DAG in Figure 3 illustrates that the prediction is directly related to previous diagnoses, age, and prescriptions, yet race could indirectly influence the prediction through its impact on access to healthcare, which in turn influences historical diagnoses [pfohl_counterfactual_2019]. When predicting length of stay in the hospital, counterfactual fairness is satisfied if, for example, the predicted length of stay for a black person is the same prediction they would have received if they were counterfactually born a white person, given all other variables in the DAG are held constant. However, counterfactual fairness can generally not be quantified from observed data [kusner_counterfactual_2017, salimi2019interventional, fawkes2022selection].
Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the prediction [pearl2009causality, mitchell_algorithmic_2021, alves_survey_2023]. Conditional counterfactual parity is a weaker criteria that considers the average causal effect of the protected attribute on the prediction conditional on a set of relevant observed variables [kusner_counterfactual_2017]. In our length of stay prediction example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the length of stay in the hospital does not change if their race were counterfactually altered. These definitions, as well as extensions of the previously introduced group and individual fairness measures, are formally introduced in Supplementary Section 6. Further work on causal fairness, such as path-specific and interventional fairness, are detailed in [chiappa_path-specific_2019, nabi_fair_2018, wu_pc-fairness_2019, zhang_causal_2017, salimi2019interventional, zuo_interventional_2024].
Figure 3: Directed acyclic graph (DAG) for predicting length of stay in the hospital. The DAG illustrates the causal relationships among various factors influencing the predicted length of hospital stay. Historical diagnoses, age, and prescriptions have a direct influence on predicted length of stay. Race indirectly affects the prediction through access to care, which in turn influences historical diagnoses.
DISCUSSION
In this tutorial, we introduced three primary notions of fairness: group, individual, and causal fairness. We discussed incompatibilities among different fairness definitions and also illustrated group fairness with real-world EHR data using our R package. We close by commenting on some ongoing research relevant to the topics introduced in this tutorial.
BIAS MITIGATION
While this paper focuses on how to define fairness, strategies for mitigating bias within ML applications are an equally active area of research. A recent systematic review presents various bias mitigation strategies and their use in practice [hort2023bias]. A necessary step in any bias mitigation strategy is to select the fairness definition that the mitigation strategy will attempt to enforce. Group fairness metrics are the most common targets, though alternative strategies have recently been proposed [goethals2024beyond]. Broadly, bias mitigation can be performed by de-biasing the training data (pre-processing), during the model training (in-processing), and after model training (post-processing). Pre-processing techniques include resampling or reweighting samples to correct imbalances among groups. One approach to in-processing is to include additional regularization terms to penalize a model’s deviation from a pre-specified fairness metric. In post-processing, for example, a technique to achieve equalized odds is to set distinct decision thresholds for different groups [hardt_equality_2016]. However, a practical consideration in any bias mitigation strategy is the well-known trade-off between achieving fairness and high overall accuracy and calibration [menon_cost_2018, brahmbhatt2023towards, mbakwe2023fairness]. Strategies for bias mitigation throughout the ML pipeline have recently been discussed in [gichoya2023ai].
REVISITING THE INCOMPATIBILITIES
The group fairness incompatibilities may be discouraging for practitioners aiming to achieve fairness. For example, the theorems show that it is not possible to equalize the group-specific true positive rates and positive predictive values when predicting an outcome that has different prevalence across groups of interest. A relaxed notion of fairness, known as approximate fairness or $\epsilon$-fairness, allows for a small margin of error in the group fairness metrics presented in Table 3. [denis_fairness_2023, bell2023possibility]. That is, rather than enforcing exact equality between metrics, approximate fairness deems a model fair if the difference between metrics does not exceed some small value, $\epsilon$. [bell2023possibility] demonstrates that it is possible for an model to satisfy approximate fairness across the false negative rate, false positive rate, and positive predictive value, even when moderate prevalence differences between groups exist. This finding offers a promising pathway for applying fairness in real-world settings where exact parity is often unachievable.
MISSING DATA
Missing data is a critical challenge to fairness evaluation in healthcare applications. Traditional imputation methods used to address missingness in features have been observed to exacerbate gaps in fairness and fairness-aware imputation strategies have been recently proposed [zhang2021fairness, zhang2022fairness]. Missingness in the outcome as well as the protected attribute can also occur, which brings challenges to both model training and evaluation. For example, in many EHR and medical imaging applications, the outcome of interest is only available for a small subset of data as it is obtained from time-consuming and expensive annotation. Semi-supervised strategies for fair model training and fairness evaluation are an active area of research [wick2019unlocking, ji2020can, choi2020fair, gao2024semisupervised].
In many settings, the protected attribute of interest is may also be unavailable or measured incorrectly. Note that some protected attributes are missing from Table 2, such as religion, gender identity, and sexual orientation, which are rarely available in healthcare data. For instance, most healthcare organizations have yet to integrate sexual orientation data collection into routine care due to the absence of standardized collection practices, the inability of EHR systems to capture structured data, and the socio-cultural challenges of self-reporting [grasso2019planning, kronk_transgender_2022, ruberg2020data]. Advances in fairness have therefore largely omitted sexual orientation, which underscores the necessity of future research to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [tomasev2021fairness].
CONCLUSION
Fairness is a framework to prevent the reinforcement or exacerbation of existing inequities when integrating ML models in healthcare. We provide an introduction to the landscape of fairness, including how bias can arise and a summary and comparison of existing metrics. We also introduced a group fairness R package for researchers to apply in their own problems.
FUNDING
J. Gronsbell is grateful for support of an NSERC Discovery Grant (RGPIN-2021-03734) and a University of Toronto Data Science Institute Seed Funding for Methodologists Grant.
AUTHOR CONTRIBUTIONS
JeG conceived and supervised the study. JeG, JiG, and BC drafted the manuscript. JiG and BC conducted experiments on simulated and real clinical data and developed the R package. PV provided clinical oversight, ZM and CH provided additional guidance on ML, and HT provided ethical oversight. All authors provided valuable feedback on the manuscript.
[…]

### v2

@@
 Table 1 details some common sources of bias together with an example from the healthcare literature.
-A more comprehensive review of causes of unfairness and applications outside of healthcare can be found in [suresh2021framework] an [mehrabi_survey_2021].
+A more comprehensive review of causes of unfairness and applications outside of healthcare can be found in [suresh2021framework] and [mehrabi_survey_2021].
 It is important to note bias should be a primary consideration throughout the ML pipeline, from problem formulation to model deployment [gichoya2023ai].
@@
 NOTATION
-We introduce a small amount of notation to introduce mathematical definitions of various criteria that complement our discussion in the main text.
+We use a small amount of notation to introduce mathematical definitions of various criteria that complement our discussion in the main text.
 We denote the outcome of interest as $Y$, the features used for model training as ${\bf X}$ [^2] Note that ${\bf X}$ does not have to be a vector; it could be an image, text, tensor, or any other type of data., and the categorical variable for the protected attribute as $A$.
@@
 SEPARATION
-Separation requires that the risk score does not depend on the protected attribute within the positive and negative classes (i.e., $S\perp\!\!\!\perp A\mid Y)$.
+Separation requires that the prediction does not depend on the protected attribute within the positive and negative classes (i.e., $D\perp\!\!\!\perp A\mid Y)$.
 This implies that, among individuals in the positive (or negative) class, the rate of making a positive (or negative) prediction is consistent across groups.
@@
 Sufficiency aims to equalize error rates among individuals with similar predictions [castelnovo_clarification_2022].
-Formally, sufficiency requires that the label does not depend on the protected attribute given the risk score ( i.e., $Y\perp\!\!\!\perp A\mid S)$.
-The risk score is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023].
+Formally, sufficiency requires that the label does not depend on the protected attribute given the prediction ( i.e., $Y\perp\!\!\!\perp A\mid D)$.
+The prediction is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023].
 Common sufficiency-based metrics focus on equalizing the positive predictive value (known as predictive parity), both the positive and negative predictive value (known as conditional use accuracy equality), and calibration (known as well-calibration )[chouldechova_fair_2017, berk_fairness_2018, kleinberg2017inherent, guo_calibration_2017].
@@
 For example, in models predicting in-hospital patient deterioration, achieving predictive parity ensures that the allocation of intensive care unit beds is based solely on medical need rather than protected attributes so that care is provided to those most in need [byrd2023validation].
-In terms of calibration-based metrics, a commercial algorithm to enroll US patients into high-risk was shown to violate well-calibration.
+In terms of calibration-based metrics, a commercial algorithm to enroll patients into a high-risk care management program was shown to violate well-calibration.
 At equivalent risk scores, Black patients were in significantly poorer health than white patients and therefore had to be sicker for enrollment into the program [obermeyer_dissecting_2019].
@@
 ation
-Independence ($S\perp\!\!\!\perp A$)
+Independence ($D\perp\!\!\!\perp A$)
 Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In all-cause mortality prediction, the probability of the model predicting mortality is the same for males and females.
 Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In all-cause mortality prediction, the probability of the model predicting mortality is the same for males and females after adjusting for age and pre-existing conditions.
-Separation ($S\perp\!\!\!\perp A\mid Y$)
+Separation ($D\perp\!\!\!\perp A\mid Y$)
 Equalized Odds $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\text{and}$
@@
 Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average risk scores among males and females without lung cancer are equal [ueda_artificial_2021].
-Sufficiency ($Y\perp\!\!\!\perp A\mid S$)
+Sufficiency ($Y\perp\!\!\!\perp A\mid D$)
 Conditional Use Accuracy Equality $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})$ and
@@
 Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that prediction is the same for Black and white patients [raza2023auditing].
-Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all classification thresholds for males and females [roosli_2022_peeking].
-Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all classification thresholds [roosli_2022_peeking].
+Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the risk score for males and females [roosli_2022_peeking].
+Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the risk score [roosli_2022_peeking].
 Other
@@
 We introduce one way of incorporating causality into what it means to be fair based on counterfactuals, which are simply “what-if” statements [pearl2009causal, hernan2010causal].
-In the context of fairness, the counterfactual is most often the unobserved prediction that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new].
-One of the first counterfactual-based fairness criteria, simply called counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables are held constant [kusner_counterfactual_2017].
+In the context of fairness, the counterfactual can be the unobserved prediction that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new].
+One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables are held constant [kusner_counterfactual_2017].
 To better explain this concept, we briefly introduce directed acyclic graphs (DAGs).
@@
 The DAG in Figure 3 illustrates that the prediction is directly related to previous diagnoses, age, and prescriptions, yet race could indirectly influence the prediction through its impact on access to healthcare, which in turn influences historical diagnoses [pfohl_counterfactual_2019].
-When predicting length of stay in the hospital, counterfactual fairness is satisfied if, for example, the predicted length of stay for a black person is the same prediction they would have received if they were counterfactually born a white person, given all other variables in the DAG are held constant.
+When predicting length of stay in the hospital, counterfactual fairness is satisfied if, for example, the predicted length of stay for a black person is the same prediction they would have received if they were counterfactually born a white person, given all other variables are held constant.
 However, counterfactual fairness can generally not be quantified from observed data [kusner_counterfactual_2017, salimi2019interventional, fawkes2022selection].
 Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the prediction [pearl2009causality, mitchell_algorithmic_2021, alves_survey_2023].
-Conditional counterfactual parity is a weaker criteria that considers the average causal effect of the protected attribute on the prediction conditional on a set of relevant observed variables [kusner_counterfactual_2017].
+Conditional counterfactual fairness is a stronger criteria that considers the total effect of the protected attribute on the prediction conditional on a set of relevant observed variables [kusner_counterfactual_2017].
 In our length of stay prediction example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the length of stay in the hospital does not change if their race were counterfactually altered.
@@
 Figure 3: Directed acyclic graph (DAG) for predicting length of stay in the hospital.
-The DAG illustrates the causal relationships among various factors influencing the predicted length of hospital stay.
+The DAG illustrates the relationships among various factors influencing the predicted length of hospital stay.
 Historical diagnoses, age, and prescriptions have a direct influence on predicted length of stay.
@@
 Group fairness metrics are the most common targets, though alternative strategies have recently been proposed [goethals2024beyond].
-Broadly, bias mitigation can be performed by de-biasing the training data (pre-processing), during the model training (in-processing), and after model training (post-processing).
+Broadly, bias mitigation can be performed by de-biasing the training data (pre-processing), during model training (in-processing), and after model training (post-processing).
 Pre-processing techniques include resampling or reweighting samples to correct imbalances among groups.
@@
 Missing data is a critical challenge to fairness evaluation in healthcare applications.
-Traditional imputation methods used to address missingness in features have been observed to exacerbate gaps in fairness and fairness-aware imputation strategies have been recently proposed [zhang2021fairness, zhang2022fairness].
+Traditional imputation methods used to address missingness in features have been observed to worsen gaps in fairness and fairness-aware imputation strategies have been recently proposed [zhang2021fairness, zhang2022fairness].
 Missingness in the outcome as well as the protected attribute can also occur, which brings challenges to both model training and evaluation.
@@
 Semi-supervised strategies for fair model training and fairness evaluation are an active area of research [wick2019unlocking, ji2020can, choi2020fair, gao2024semisupervised].
-In many settings, the protected attribute of interest is may also be unavailable or measured incorrectly.
+In many settings, the protected attribute of interest may also be unavailable or measured incorrectly.
 Note that some protected attributes are missing from Table 2, such as religion, gender identity, and sexual orientation, which are rarely available in healthcare data.